# How do Laws Produce the Future?

### v2

@@ 1 Introduction
The dynamic production account of physical laws can be contrasted with two alternatives. [^1] There are other options. Adlam, (2022a), Chen & Goldstein, (2022), and Chen, (2024) discuss the universals account (Platonic reductionism), the powers account (Aristotelian reductionism), and the modal structure account. The first is the Humean Mill-Ramsey-Lewis best systems account, according to which the laws do not truly govern what happens in nature, but instead merely offer concise and informative descriptions of all the things that happen throughout the entire history of the universe (including its full past and future). [^2] See Loewer, (1996); Lange, (2008). Put poetically, the laws are patterns in the mosaic that is our universe. The second is a recent alternative, put forward by Chen & Goldstein, (2022), Adlam, (2022a), and Meacham, (2022, forthcoming), according to which the laws govern without dynamic production by placing constraints on possible histories for the universe (constraints that may or may not look like laws of time evolution).
@@ 2 The Time Evolution Paradigm
as in Lange, (2005, sec. 2). This velocity is depicted in figure 1. The future velocity, by contrast, would be defined by considering arbitrarily small time intervals to the future. In ordinary circumstances, the past and future velocities will agree. However, they could in principle come apart if there is a kink in the trajectory of a body through space and time. Whether past and future velocities are permitted to come apart will depend on the laws of the theory at hand.
@@
This formulation allows for non-dynamical laws in addition to dynamical laws. As a candidate example of a non-dynamical law, Maudlin, (2007, pg. 13) gives Newton’s law of gravitation, which uses present positions and masses to specify the present force on a given body. Maudlin describes this law as an “adjunct principle” that allows the dynamical law $\vec{F}=m\vec{a}$ to predict future time evolution. Chen & Goldstein, (2022, pg. 60) consider the example of Gauss’s law, $\vec{\nabla}\cdot\vec{E}=4\pi\rho$, which can be regarded as a non-dynamical law relating the present divergence of the electric field to the present distribution of charge. The requirement in the second condition that non-dynamical laws “express relations between present properties or past neighborhood properties” rules out certain putative non-dynamical laws, such as those that posit probability distributions over initial conditions (Chen & Goldstein,, 2022, sec. 3.3.3). [^10] One could still posit a probability distribution over initial conditions to explain thermodynamic asymmetries, as Albert, (2000) does with his Past Hypothesis and Statistical Postulate. Within the time evolution paradigm, such a probability distribution would have to be regarded as a hypothesis (or postulate), not a law of nature. The third condition, that the laws be either deterministic or stochastic, is meant to exclude laws that give incomplete stories about the future: neither determining the future uniquely nor giving definite probabilities for different futures. As an example of such an incomplete theory, consider a version of electromagnetism where the vector and scalar potentials are taken to be the proper way to represent the electromagnetic field and no gauge fixing condition is imposed. Then, multiple future evolutions of the potentials will be compatible with the past (Maudlin,, 2018, pg. 14).
@@
The third condition of the time evolution paradigm requires the laws to be either deterministic or stochastic. This is a natural requirement for the dynamic production account and it is one that is imposed by Maudlin, (2007, pg. 14–20, 174–183). To dynamically produce the future, the laws must be sufficiently specific that they either yield a unique future or a probability distribution over possible futures. In a deterministic theory, the dynamical laws must together say enough about past-to-future neighborhood properties to specify a single future evolution—when applied at all times, alongside the non-dynamical laws. This is the simplest way that the laws might produce the future and a narrow version of the dynamic production account might require the laws to take this form. By contrast, stochastic laws allow for multiple futures and give probabilities for each. [^15] These probabilities should be understood as basic propensities that are not to be reduced to anything more fundamental, such as long-run frequencies. See Maudlin, (2007, pg. 17–20); Frigg & Hoefer, (2007); Loewer, (2012, pg. 118). Such laws can arguably also be regarded as producing the future, though they do so in a fundamentally random way where different futures have different chances of occurring. One can write down laws that are neither deterministic nor stochastic, deeming multiple futures to be compatible with a given past. It is hard to see how such laws could be understood as producing the future. Such laws do not specify, even in a probabilistic way, which future is to be produced (leaving open multiple possibilities). Here are three case where that might occur: First, as was mentioned previously, this kind of indeterminism arises if electromagnetism is formulated in terms of potentials without a gauge fixing condition. Maudlin, (2018) advocates adopting a gauge-fixing condition to avoid this indeterminism. Second, this kind of indeterminism may also occur in general relativity if one allows for the possibility of closed time-like curves (Earman et al.,, 2024). Maudlin, (2007, pg. 175, 189–191) forbids closed time-like curves as incompatible with his picture of dynamic production. Third, there have been worries raised about Newton’s laws of classical mechanics leaving open multiple futures and thus not being as deterministic as they appear (Earman,, 2004; Wilson,, 2009; van Strien,, 2021). A defender of dynamic production would seek sets of laws that avoid such indeterminism.
@@
A defender of dynamic production may allow that it is metaphysically possible to have laws of nature that violate the time evolution paradigm and cannot be understood as dynamically producing the future from the past. Adopting this perspective, dynamic production is not part of what it takes for something to be a law. It is merely a feature of our laws. It is difficult to say what it takes in general for something to be a law, but you might say that at a minimum it must constrain what is physically possible. Here the dynamic production account can make peace with constraint accounts (Chen & Goldstein,, 2022; Adlam,, 2022a; Meacham,, 2022, forthcoming). Proponents of constraint accounts prefer the flexibility of their picture, arguing that the laws of our world may well turn out not to include dynamical laws (or to otherwise violate the time evolution paradigm) and that our philosophical theorizing should not preclude this possibility. [^16] In addition to Chen & Goldstein, (2022); Adlam, (2022a); Meacham, (2022, forthcoming), see also Wharton, (2015); Lam & Wüthrich, (2023). Proponents of dynamic production stick their necks out, betting that the final laws will fit the time evolution paradigm because our most successful extant theories are theories of time evolution (or at least can be formulated as theories of time evolution). This debate hinges primarily on difficult questions about the interpretation of general relativity and quantum field theory that will only be discussed briefly in section 7. Here we focus on theories that seem like they should fit the time evolution paradigm, to better understand how dynamic production works in these cases.
@@ 6.1 The evolution of the electromagnetic field
where we will again treat $f$ as a fixed source function defined across space and time. The wave equations for the potentials, (17) and (18), fit this general form. Writing this wave equation as a spatiotemporally local dynamical law is not as straightforward as you might expect, but here is one way to do it. Let us focus our attention on the evolution at the origin, noting that any point in space can be treated as the origin. At the origin, [^34] Away from the origin, the Laplacian cannot be put in this form and must include derivatives with respect to $\theta$ and $\phi$ in addition to $r$ (as is evident in the standard expression for the Laplacian in spherical coordinates). the Laplacian $\nabla^{2}=\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}+\frac{\partial^{2}}{\partial z^{2}}$ can be rewritten in spherical coordinates as an integral over contributions from second derivatives with respect to the radial coordinate $r$ for different choices of $\theta$ (the azimuthal angle) and $\phi$ (the polar angle)—seeing how $u$ changes as you look out in different directions from the origin: [^35] To derive (33), plug the expansion $\frac{\partial}{\partial r}=\cos\theta\sin\phi\frac{\partial}{\partial x}+\sin\theta\sin\phi\frac{\partial}{\partial y}+\cos\phi\frac{\partial}{\partial z}$ (31) into the integral over the solid angle. The cross terms drop out and you are left with $\int\frac{d\Omega}{4\pi}\left(3\frac{\partial^{2}}{\partial r^{2}}\right)u$ $=\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}+\frac{\partial^{2}}{\partial z^{2}}\right)u=\nabla^{2}u\ .$ (32)
@@ 6.4 The whole package
We have seen that this formulation of electromagnetism with point charges satisfies the first two conditions of the relativistic time evolution paradigm: the dynamical and non-dynamical laws take the correct form. We have not yet settled whether the formulation satisfies the third condition as we have not yet settled whether the theory is locally deterministic. To do so, we must figure out whether, in general, an arbitrarily thin law-abiding time slice of a contracting light-cone yields a unique remainder of that light-cone. This can be proven for the wave equations governing the evolution of the fields, [^43] See the references in footnote 31. but proving it for the entire theory requires looking at the coupled equations for the field and particle dynamics. [^44] See Frisch, (2004, sec. 2); Frisch, (2005, pg. 32–35). Doing this honestly would require a solution to the aforementioned problems of self-interaction. There is another complication as well. Hartenstein & Hubert, (2021) have shown that specifying law-abiding combinations of electromagnetic field and point charge states at an instant generically yields infinities or discontinuities that propagate along the future light-cone (called “shock fronts”) and lead to ill-defined dynamics. [^45] See also Lazarovici, (2018, sec. 8.1). These shock fronts occur because the arbitrarily chosen field around a point charge does not capture the way that the charge’s past motion would have acted as a source for the field. For example, at some moment you might have a particle with an initial velocity surrounded by the Coulomb field of a stationary point charge. That is not compatible with any reasonable past. Our formulation of determinism (from section 2) requires that we start with an arbitrarily short time interval where the laws are obeyed and that starting point should help in addressing these problems. We are assuming that the interaction between particle and field has been well-behaved in the immediate past and then asking whether the future evolution is well-behaved and unique.
@@
Shifting now to consider electromagnetism with a continuous charge distribution, the wave equations for the potentials, (42) and (43), would again be spatiotemporally local dynamical laws. In addition, we would need law(s) determining the evolution of the charged matter. The Lorentz force density law (48) can be formulated as an acceptable non-dynamical law, but it will not be a fundamental law for certain classical completions of electromagnetism. In particular, if we model the evolution of matter using the Klein-Gordon or Dirac equations in Maxwell-Klein-Gordon or Maxwell-Dirac field theory, it is not needed. As was discussed earlier, the Klein-Gordon equation can be formulated as a spatiotemporally local dynamical law and one might find a way to do so for the Dirac equation. The Maxwell-Klein-Gordon and Maxwell-Dirac field theories could then satisfy the first two conditions of the relativistic time evolution paradigm. They should satisfy the third condition as well—the free Klein-Gordon and Dirac equations are locally deterministic [^46] See Wald, (1984, sec. 10.1); Thaller, (1992, sec. 1.5); Strauss, (2008, pg. 234, problem 8) and the introduction of interactions does not bring about any infinite or ill-defined self-interaction. [^47] See Sebens, (2023).
@@ 7 The Path Ahead
To build up to quantum field theory, let us start by discussing the non-relativistic quantum mechanics of a fixed number of particles. In that context, there exist a handful of “interpretations of quantum mechanics” that give explicit proposals about the laws and ontology—such as GRW theory, Bohmian mechanics, the many-worlds interpretation, and the many interacting worlds approach. These four options can all be fit straightforwardly into the time evolution paradigm, though there are other “retrocausal,” “all-at-once,” or “non-Markovian” options [^48] See Wharton, (2014); Adlam, (2018, 2022a, 2022b, 2023); Friederich & Evans, (2023); Barandes, (2023). that cannot. Focusing on the four aforementioned options, the dynamics will be either first-order or second-order depending on the interpretation. The many-worlds interpretation simply takes the first-order Schrödinger equation to give the dynamics of the wave function. This is naturally interpreted as taking the wave function at a moment as input (with no need to look to the arbitrarily-short past) and returning as output the way the wave function changes into the arbitrarily short future (Easwaran,, 2014, pg. 857). GRW theory introduces stochastic exceptions to the Schrödinger equation that can be understood as giving probabilities for different behavior in the arbitrarily short future (Sebens,, 2015a, sec. 2). Bohmian mechanics is often presented as adding a first-order guidance equation, that is another law beyond the Schrödinger equation, determining the velocities of particles from their locations and the wave function (velocities that could be interpreted as future neighborhood properties). However, a second-order guidance equation can be used instead provided we impose the first-order guidance equation as a restriction on initial conditions (Goldstein & Struyve,, 2015; Dewdney,, 2023). In the many interacting worlds approach, the dynamics is second-order as particles are reacting to forces by Newton’s second law and (at the fundamental level) there is no wave function evolving by the Schrödinger equation (Hall et al.,, 2014; Sebens,, 2015b; Ghadimi et al.,, 2018). The fact that quantum theories normally posit first-order dynamics has been used by Builes & Impagnatiello, (forthcoming) to argue that we live in a world where the present moment is enough to produce the future (the arbitrarily-short past is not needed as input)—in their words, “our universe is Markovian.” I prefer to leave open whether the fundamental dynamics will turn out to be first-order or second-order.

### v3

@@ 1 Introduction
The dynamic production account of physical laws can be contrasted with two alternatives. [^1] There are other options. Adlam, (2022a), Chen & Goldstein, (2022), and Chen, (2024) discuss the universals account (Platonic reductionism), the powers account (Aristotelian reductionism), and the modal structure account. The first is the Humean Mill-Ramsey-Lewis best systems account, according to which the laws do not truly govern what happens in nature, but instead merely offer concise and informative descriptions of all the things that happen throughout the entire history of the universe (including its full past and future). [^2] See Loewer, (1996); Lange, (2008). Put poetically, the laws are patterns in the mosaic that is our universe. The second is a recent alternative, put forward by Adlam, (2022a), Chen & Goldstein, (2022), and Meacham, (2022, forthcoming), according to which the laws govern without dynamic production by placing constraints on possible histories for the universe (constraints that may or may not look like laws of time evolution).
@@ 2 The Time Evolution Paradigm
as in Lange, (2005, sec. 2). This velocity is depicted in figure 1. The future velocity, by contrast, would be defined by considering arbitrarily small time intervals to the future. In ordinary circumstances, the past and future velocities will agree. However, they could, in principle, come apart if there is a kink in the trajectory of a body through space and time. Whether past and future velocities are permitted to come apart will depend on the laws of the theory at hand.
@@
This formulation allows for non-dynamical laws in addition to dynamical laws. As a candidate example of a non-dynamical law, Maudlin, (2007, pg. 13) gives Newton’s law of gravitation, which uses present positions and masses to specify the present force on a given body. Maudlin describes this law as an “adjunct principle” that allows the dynamical law $\vec{F}=m\vec{a}$ to predict future time evolution. Chen & Goldstein, (2022, pg. 60) consider the example of Gauss’s law, $\vec{\nabla}\cdot\vec{E}=4\pi\rho$, which can be regarded as a non-dynamical law relating the present divergence of the electric field to the present distribution of charge. The requirement in the second condition that non-dynamical laws “express relations between present properties or past neighborhood properties” rules out certain putative non-dynamical laws, such as those that posit probability distributions over initial conditions (Chen & Goldstein,, 2022, sec. 3.3.3). [^10] One could still posit a probability distribution over initial conditions to explain thermodynamic asymmetries, as Albert, (2000) does with his Past Hypothesis and Statistical Postulate. Within the time evolution paradigm, such a probability distribution would have to be regarded as a hypothesis (or postulate), not a law of nature. The third condition, that the laws be either deterministic or stochastic, is meant to exclude laws that give incomplete stories about the future: neither determining the future uniquely nor giving a definite probability distribution over different futures. As an example of such an incomplete theory, consider a version of electromagnetism where the vector and scalar potentials are taken to be the proper way to represent the electromagnetic field and no gauge fixing condition is imposed. Then, multiple future evolutions of the potentials will be compatible with the past (Maudlin,, 2018, pg. 14).
@@
The third condition of the time evolution paradigm requires the laws to be either deterministic or stochastic. This is a natural requirement for the dynamic production account and it is one that is imposed by Maudlin, (2007, pg. 14–20, 174–183). To dynamically produce the future, the laws must be sufficiently specific that they either yield a unique future or a probability distribution over possible futures. In a deterministic theory, the dynamical laws must together say enough about past-to-future neighborhood properties to specify a single future evolution—when applied at all times, alongside the non-dynamical laws. This is the simplest way that the laws might produce the future and a narrow version of the dynamic production account might require the laws to take this form. By contrast, stochastic laws allow for multiple futures and provide a probability distribution over such futures. [^15] In the dynamic production account, the probabilities that appear in stochastic laws should be understood as basic propensities that are not to be reduced to anything more fundamental (such as long-run frequencies). See Maudlin, (2007, pg. 17–20); Frigg & Hoefer, (2007); Loewer, (2012, pg. 118). Such laws can arguably also be regarded as producing the future, though they do so in a fundamentally random way. One can write down laws that are neither deterministic nor stochastic, deeming multiple futures to be compatible with a given past. It is hard to see how such laws could be understood as producing the future. Such laws do not specify, even in a probabilistic way, which future is to be produced (leaving open multiple possibilities). Here are three cases where that might occur: First, as was mentioned previously, this kind of indeterminism arises if electromagnetism is formulated in terms of potentials without a gauge fixing condition. Maudlin, (2018) advocates adopting a gauge-fixing condition to avoid this indeterminism. Second, this kind of indeterminism may also occur in general relativity if one allows for the possibility of closed time-like curves (Earman et al.,, 2024). Maudlin, (2007, pg. 175, 189–191) forbids closed time-like curves as incompatible with his picture of dynamic production. Third, there have been worries raised about Newton’s laws of classical mechanics leaving open multiple futures and thus not being as deterministic as they appear (Earman,, 2004; van Strien,, 2021; Wilson,, 2009). A defender of dynamic production would seek sets of laws that avoid such indeterminism.
@@
A defender of dynamic production may allow that it is metaphysically possible to have laws of nature that violate the time evolution paradigm and cannot be understood as dynamically producing the future from the past. Adopting this perspective, dynamic production is not part of what it takes for something to be a law. It is merely a feature of our laws. It is difficult to say what it takes in general for something to be a law, but you might say that at a minimum it must constrain what is physically possible. Here the dynamic production account can make peace with constraint accounts (Adlam,, 2022a; Chen & Goldstein,, 2022; Meacham,, 2022, forthcoming). Proponents of constraint accounts prefer the flexibility of their picture, arguing that the laws of our world may well turn out not to include dynamical laws (or to otherwise violate the time evolution paradigm) and that our philosophical theorizing should not preclude this possibility. [^16] In addition to Chen & Goldstein, (2022); Adlam, (2022a); Meacham, (2022, forthcoming), see also Wharton, (2015); Lam & Wüthrich, (2023). Proponents of dynamic production stick their necks out, betting that the final laws will fit the time evolution paradigm because our most successful extant theories are theories of time evolution (or at least can be formulated as theories of time evolution). This debate hinges primarily on difficult questions about the interpretation of general relativity and quantum field theory that will only be discussed briefly in section 7. Here we focus on theories that seem like they should fit the time evolution paradigm, to better understand how dynamic production works in these cases.
@@ 6.1 The evolution of the electromagnetic field
where we will again treat $f$ as a fixed source function defined across space and time. The wave equations for the potentials, (17) and (18), fit this general form. Writing this wave equation as a spatiotemporally local dynamical law is not as straightforward as you might expect, but here is one way to do it. Let us focus our attention on the evolution at the origin, noting that any point in space can be treated as the origin. At the origin, [^34] Away from the origin, the Laplacian cannot be put in this form and must include derivatives with respect to $\theta$ and $\phi$ in addition to $r$ (as is evident in the standard expression for the Laplacian in spherical coordinates). the Laplacian $\nabla^{2}=\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}+\frac{\partial^{2}}{\partial z^{2}}$ can be rewritten in spherical coordinates as an integral over contributions from second derivatives with respect to the radial coordinate $r$ for different choices of $\theta$ (the azimuthal angle) and $\phi$ (the polar angle)—seeing how $u$ changes as you look out in different directions from the origin: [^35] I have not seen (33) presented elsewhere. To derive (33), plug the expression for the radial derivative given a particular choice of $\theta$ and $\phi$, $\frac{\partial}{\partial r}=\cos\theta\sin\phi\frac{\partial}{\partial x}+\sin\theta\sin\phi\frac{\partial}{\partial y}+\cos\phi\frac{\partial}{\partial z}\ ,$ (31) into the integral over these angles on the right-hand side of (33). When you perform the integrals over $\theta$ and $\phi$, the cross terms drop out and you are left with the Laplacian (at the origin): $3\int\frac{d\theta d\phi}{4\pi}\sin\phi\;\frac{\partial^{2}}{\partial r^{2}}u$ $=3\int\frac{d\theta d\phi}{4\pi}\bigg{(}\cos^{2}\theta\sin^{3}\phi\frac{\partial^{2}}{\partial x^{2}}+\sin^{2}\theta\sin^{3}\phi\frac{\partial^{2}}{\partial y^{2}}+\cos^{2}\phi\sin\phi\frac{\partial^{2}}{\partial z^{2}}$ $\quad+2\cos\theta\sin\theta\sin^{3}\phi\frac{\partial}{\partial x}\frac{\partial}{\partial y}+2\sin\theta\sin^{2}\phi\cos\phi\frac{\partial}{\partial y}\frac{\partial}{\partial z}+2\cos\theta\sin^{2}\phi\cos\phi\frac{\partial}{\partial x}\frac{\partial}{\partial z}\bigg{)}u$ $=\left(\frac{\partial^{2}}{\partial x^{2}}+\frac{\partial^{2}}{\partial y^{2}}+\frac{\partial^{2}}{\partial z^{2}}\right)u=\nabla^{2}u\ .$ (32)
@@ 6.4 The whole package
We have seen that this formulation of electromagnetism with point charges satisfies the first two conditions of the relativistic time evolution paradigm: the dynamical and non-dynamical laws take the correct form. We have not yet settled whether the formulation satisfies the third condition as we have not yet settled whether the theory is locally deterministic. To do so, we must figure out whether, in general, an arbitrarily thin law-abiding time slice of a contracting light-cone yields a unique remainder of that light-cone. This can be proven for the wave equations governing the evolution of the fields, [^43] See the references in footnote 31. but proving it for the entire theory requires looking at the coupled equations for the field and particle dynamics. [^44] See Frisch, (2004, sec. 2); Frisch, (2005, pg. 32–35). Doing this honestly would require a solution to the aforementioned problems of self-interaction. There is another complication as well. Hartenstein & Hubert, (2021) have shown that specifying law-abiding combinations of electromagnetic field and point charge states at an instant generically yields infinities or discontinuities that propagate along the future light-cone (called “shock fronts”) and leads to ill-defined dynamics. [^45] See also Lazarovici, (2018, sec. 8.1). These shock fronts occur because the arbitrarily chosen field around a point charge does not capture the way that the charge’s past motion would have acted as a source for the field. For example, at some moment you might have a particle with an initial velocity surrounded by the Coulomb field of a stationary point charge. That is not compatible with any reasonable past. Our formulation of determinism (from section 2) requires that we start with an arbitrarily short time interval where the laws are obeyed and that starting point should help in addressing these problems. We are assuming that the interaction between particle and field has been well-behaved in the immediate past and then asking whether the future evolution is well-behaved and unique.
@@
Shifting now to consider electromagnetism with a continuous charge distribution, the wave equations for the potentials, (42) and (43), would again be spatiotemporally local dynamical laws. In addition, we would need law(s) determining the evolution of the charged matter. The Lorentz force density law (48) can be formulated as an acceptable non-dynamical law, but it will not be a fundamental law for certain classical completions of electromagnetism. In particular, if we model the evolution of matter using the Klein-Gordon or Dirac equations in Maxwell-Klein-Gordon or Maxwell-Dirac field theory, it is not needed. As was discussed earlier, the Klein-Gordon equation can be formulated as a spatiotemporally local dynamical law and one might find a way to do so for the Dirac equation. The Maxwell-Klein-Gordon and Maxwell-Dirac field theories could then satisfy the first two conditions of the relativistic time evolution paradigm. They should satisfy the third condition as well—the free Klein-Gordon and Dirac equations are locally deterministic [^46] See Wald, (1984, sec. 10.1); Thaller, (1992, sec. 1.5); Strauss, (2008, pg. 234, problem 8). and the introduction of interactions does not bring about any infinite or ill-defined self-interaction. [^47] See Sebens, (2023).
@@ 7 The Path Ahead
To build up to quantum field theory, let us start by discussing the non-relativistic quantum mechanics of a fixed number of particles. In that context, there exist a handful of “interpretations of quantum mechanics” that give explicit proposals about the laws and ontology—such as GRW theory, Bohmian mechanics, the many-worlds interpretation, and the many interacting worlds approach. These four options can all be fit straightforwardly into the time evolution paradigm, though there are other “retrocausal,” “all-at-once,” or “non-Markovian” options [^48] See Wharton, (2014); Adlam, (2018, 2022a, 2022b, 2023); Friederich & Evans, (2023); Barandes, (2023). that cannot. Focusing on the four aforementioned options, the dynamics will be either first-order or second-order depending on the interpretation. The many-worlds interpretation simply takes the first-order Schrödinger equation to give the dynamics of the wave function. This is naturally interpreted as taking the wave function at a moment as input (with no need to look to the arbitrarily-short past) and returning as output the way the wave function changes into the arbitrarily-short future (Easwaran,, 2014, pg. 857). GRW theory introduces stochastic exceptions to the Schrödinger equation that can be understood as giving probabilities for different behaviors in the arbitrarily-short future (Sebens,, 2015a, sec. 2). Bohmian mechanics is often presented as adding a first-order guidance equation, that is another law beyond the Schrödinger equation, determining the velocities of particles from their locations and the wave function (velocities that could be interpreted as future neighborhood properties). However, a second-order guidance equation can be used instead provided we impose the first-order guidance equation as a restriction on initial conditions (Goldstein & Struyve,, 2015; Dewdney,, 2023). In the many interacting worlds approach, the dynamics is second-order as particles are reacting to forces by Newton’s second law and (at the fundamental level) there is no wave function evolving by the Schrödinger equation (Hall et al.,, 2014; Sebens,, 2015b; Ghadimi et al.,, 2018). The fact that quantum theories normally posit first-order dynamics has been used by Builes & Impagnatiello, (forthcoming) to argue that we live in a world where the present moment is enough to produce the future (the arbitrarily-short past is not needed as input)—in their words, “our universe is Markovian.” I prefer to leave open whether the fundamental dynamics will turn out to be first-order or second-order.